OpenStudy (anonymous):

need help could someone walk me through this....6x+9-4x=3x+11

OpenStudy (anonymous):

ok. So first pick what side you want to have the x's in. normally they are on the left side.

OpenStudy (anonymous):

so get your x's on the left side following your oder of operations.

OpenStudy (anonymous):

combine the terms right

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

You should get 6x-4x-3x=11-9

OpenStudy (anonymous):

did you get this?

OpenStudy (anonymous):

rt sorry typing sucks

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

o.k now combine the terms.

OpenStudy (anonymous):

just add/ subtract respectively

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

13x-11-9

OpenStudy (anonymous):

13x=11-9

OpenStudy (anonymous):

remember we want to keep the x's with the x's and the integers with the integers.

OpenStudy (anonymous):

dont forget about your signs. Look at the signs: 6x-4x-3x

OpenStudy (anonymous):

notice that we are adding two negative numbers.

OpenStudy (anonymous):

remember (-3)+(-3)=-6

OpenStudy (anonymous):

-6-6=-12

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

the only time when two negative numbers become positive is when we multiply them

OpenStudy (anonymous):

right gotthat

OpenStudy (anonymous):

so what would -4x-3x be?

OpenStudy (anonymous):

-7x

OpenStudy (anonymous):

cool. So what is 6x-7x

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

almost. -x

OpenStudy (anonymous):

forgot 1x is same asx

OpenStudy (anonymous):

so we got the left side taken care. The right side should give us 2

OpenStudy (anonymous):

so we have -x=2. we are almost done

OpenStudy (anonymous):

all that is left is getting rid of the negative. All we need to do is divide by a negative.

OpenStudy (anonymous):

so the answer is x=-2.

OpenStudy (anonymous):

see i get lost on when to change the operation like sub to add or change neg to pos

OpenStudy (anonymous):

whenever you have the equation in just x's on one side and integers on the right, then you can divide.

OpenStudy (anonymous):

3x=53, x=53/3. Its ok to divide. 3x=53+5x, you can divide but it will just be messy: x=(53+5x)/3. Besides we have to solve for x and that just complicated things.

OpenStudy (anonymous):

3x-5x=53, -2x=53, x=-53/2

OpenStudy (anonymous):

ok so when i see a - sign its the same as neg

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

3x-5x is -2x......but when i have an equation here let me show u if u have time?

OpenStudy (anonymous):

yeah. its cool. show away.

OpenStudy (anonymous):

4(3x-5)=-4x+27

OpenStudy (anonymous):

what screws me up is when im not sure if i add or sub

OpenStudy (anonymous):

so its 12x-20=4x+27

OpenStudy (anonymous):

then move the smaller x ?

OpenStudy (anonymous):

its really up to you. It does not matter. The answer will work out the same.

OpenStudy (anonymous):

you just have to remember to keep your x's with your x's on one side of the equation and your integers on the other side.

OpenStudy (anonymous):

whether you want to keep the x on the right or left side it does not matter.

OpenStudy (anonymous):

12x-4x=27+20 -20-27=4x-12x 8x=47 -47=-8x x=47/8 47/8=x (negatives cancel.)

OpenStudy (anonymous):

makes sense sry im really not sow....lol always had problem with math

OpenStudy (anonymous):

slow..lol

OpenStudy (anonymous):

math just takes time. After graduating from high school i did my own thing for 4 year not worrying about school. After 4 yrs I decided to go to college. Took a placement test and place in pre-algebra. I have been working my way up ever since.

OpenStudy (anonymous):

just be patient with it.

OpenStudy (anonymous):

thats exactly what im doin now...but its been 14 yrs since ive been to school

OpenStudy (anonymous):

what math u doin now

OpenStudy (anonymous):

im in differential equations and linear algebra.

OpenStudy (anonymous):

cool hey thx for ur help im keith btw

OpenStudy (anonymous):

r u on here often

OpenStudy (anonymous):

no worries. yeah. Im usually asking questions as well. We all need help.

OpenStudy (anonymous):

well good luck on your math studies.

OpenStudy (anonymous):

u too take care